Use the definition of the cotangent function and the quotient rule to prove if f(x)=cotx, than f'(x)= -cosec2x.


1
Expert's answer
2021-07-12T14:04:45-0400
"f'(x)=(\cot x)'=(\dfrac{\cos x}{\sin x})'"

"=\dfrac{(\cos x)'\sin x-\cos x(\sin x)'}{\sin^2x}=\dfrac{-\sin^2x-\cos^2x}{\sin^2x}"

"=\dfrac{-(\sin^2x+\cos^2x)}{\sin^2x}=\dfrac{-1}{\sin^2x}=-\cosec^2x"